OpenStudy (anonymous):

Factor 81b^6-3d^6=I got (3(27b^6-d^6) what is the correct form to write the answer.

OpenStudy (anonymous):

do you mean to have the equals sign in there?

OpenStudy (anonymous):

not I have the answer but I don't have it in the right written form

OpenStudy (anonymous):

I'm pretty sure that you can treat 27b^6-d^6 as a difference of squares

OpenStudy (anonymous):

Your answer is a good first step, but now you need to apply the sum of two cubes formula I gave you in the last problem.

OpenStudy (anonymous):

a3 + b3 = (a + b)(a2 – ab + b2) In this case, a = 27(b^2), and b=-d^2.

OpenStudy (anonymous):

I need to buy the algebra for dummies book :)

OpenStudy (anonymous):

completely ignore what i said, totally wrong approach

OpenStudy (anonymous):

Sorry, a = 3(b^2)

OpenStudy (anonymous):

so it will be (b+3)(b^3-27b-d6)?

OpenStudy (anonymous):

Nah, it'd be 3(3(b^2)-d^2)(9b^4 + 3b^2d^2 - d^4).

OpenStudy (anonymous):

oh wow I was way off.